# Investigating Feature and Model Importance in Android Malware Detection: An Implemented Survey and Experimental Comparison of ML-Based Methods

### v3

@@ 7. Ensemble Models
8. Discussion
@@
This study allowed us to take a broader perspective on the Android malware-detection literature and, arguably for the first time, see how the different ML-based approaches to solving this problem work within contemporary Android systems. Some of the observations were surprising. For instance, our results suggest that there is generally little advantage to using dynamic analysis techniques, and that very accurate models can be built using relatively simple static features, particularly API call usage. The only scenario where dynamic analysis seemed to bring an advantage was when we built models based around network activity. However, this is expensive to implement in practice, since collection of network activity requires dynamic analysis at both training and inference time, and relies on features which are both difficult to collect and potentially brittle.
@@
A recurring observation was the challenge caused by the growing complexity of Android. For instance, we observed that the number of API calls has grown by two orders of magnitude since early work on ML-based Android malware detection. This means the use of feature selection is now very important in order to reduce feature numbers down to manageable levels. Within this context, an important observation is that feature selection does not just reduce the computational resources needed for training; it also, in most cases, increases the accuracy of the resulting models. Most notably, we have shown that reducing the number of API calls by 95% leads to higher levels of malware discrimination. In general, we found mutual information and chi-square to be the most effective feature selectors, and did not observe any benefit to using a more complex model-based feature selection approach. The Drebin feature set has historically played an important role in ML-based malware detection studies. Although this feature set still performs reasonably well, our results suggest that it is no longer the best feature set to use, and is difficult to deploy due to the large number of features it entails when used within a contemporary Android environment. Unlike other feature sets we studied, it does not benefit from feature selection.
@@
Beyond features, we also found that model choice is important, and random forests were surprisingly consistent in being the best performing models across the majority of experiments and feature sets. SVMs also performed well in general, but were more expensive to train than random forests. Naïve Bayes models generally performed poorly, suggesting that, for malware detection, interactions between features are important. A somewhat surprising observation, given their increasing prevalence within ML more generally, is that deep neural network models did not perform particularly well. However, this may be due to the use of pre-extracted features in the majority of malware detection studies. In situations where raw data is used (e.g. raw source code), there may be benefits to using more complex deep learning models such as transformers. However, there are also significant risks of overfitting and greater sensitivity to distribution shifts (Lones, 2021).
Finally, we observed that a number of the malware detection approaches reported in the literature performed significantly less well when we reimplemented and evaluated them on a large contemporary dataset. In many cases, this may be attributed to the increasing complexity of the Android environment; however, poor experimental practice is also likely to be a contributing factor, and we have observed many instances of authors falling into common ML pitfalls (Lones, 2021), including data leakage and the use of inappropriate metrics. Either way, it shows the importance of periodically reimplementing and reevaluating studies.
9. Limitations and Future Work
The dynamic nature of the Android operating system results in new studies being published at a rapid pace. Notably, transformers are becoming increasingly popular across applied ML, and there are examples of these being applied to Android malware detection (Liu et al., 2022). However, the complexity of these models does introduce new challenges (Lones, 2021).
@@ 9. Limitations and Future Work
Our aim in this study is to sample and reimplement representative studies from the literature. It is not feasible to reimplement all existing studies, and there is inevitably some sampling bias within our approach. In particular, we focus on studies which used similar underlying features, since this enables a more meaningful basis for comparison across a standardised data set. This approach was motivated by our earlier review (Muzaffar et al., 2022) of the most common ML-based modelling approaches in Android malware detection.
It is worth noting in particular that transformers are becoming increasingly popular across applied ML, and there are examples of these being applied to Android malware detection (Oak et al., 2019; Fan et al., 2021; Saracino and Simoni, 2023; Yang et al., 2024). Unlike the approaches surveyed in this work, these approaches tend to use lower-level features, which may be constructed directly from raw data. This introduces benefits as well as challenges (Lones, 2021). An example of a transformer-based approach is the work of (Fan et al., 2021), which addresses the challenge of integrating spatial and temporal information by combining two different transformer models: one which focuses on spatial patterns and another which focuses on temporal patterns. This modelling approach also allows the system to integrate behavioural information beyond the other studies surveyed here.
There are other (non-transformer-based) studies which use broader information sources. For example, Hou et al. (Hou et al., 2017), in their HinDroid framework, consider the relationships between applications and APIs. Similarly, Ye et al. (Ye et al., 2019) utilized features that incorporated various forms of metadata from apps and devices, including IMEI numbers.
In future work, it would be interesting to implement and compare studies that incorporate these more heterogeneous features, including recent transformer-based approaches.
10. Conclusions
In this work, we have have experimentally investigated the factors that affect model accuracy when applying machine learning to Android malware detection. We focused, in particular, on the importance of feature and model selection. An important component of our investigation was the reimplementation and reevaluation of existing approaches from the literature. This allowed us both to carry out a rigorous experimental comparison of these approaches, and to determine whether they remain relevant to contemporary Android systems. We also carried out a number of new experiments to fill in the knowledge gaps found in previous studies. These experiments were supported by the collection of a large up-to-date data set comprising 124,000 Android applications, equally divided between malware and benign applications. The tools used to create this dataset, as well as the tools we developed to extract features, are provided as a resource for the community.
@@ 10. Conclusions
A key observation is that features extracted using static analysis of Android APKs are often more effective than those extracted using dynamic analysis of running applications. This is important to know, since static analysis is much cheaper than dynamic analysis. Within this context, we found that simple representations of API usage led to very effective models, with accuracies of up to 96.8%. The exception to this is when features extracted from network traffic are used to build models. In our study, it is these models that provided the highest levels of malware discrimination, achieving accuracies of up to 97.7%. However, they are expensive to train and use, and rely on features that are difficult to collect and potentially brittle. Fortunately, we also showed that ensemble models built from non-network features can achieve a similar level of discrimination (97.8%), offering a more practical alternative for real world deployment.
Other notable observations include the importance of using feature selection algorithms when dealing with the complexity of modern Android systems, the high accuracy of random forest models in comparison to other ML models (including deep learners), and the relatively poor performance of the historically-important Drebin feature set. In future work, we plan to focus on online ML-based Android malware detection. Keeping a model up to date is important for accurate malware detection, and online learning can also be useful for drift detection. The sheer number of applications released everyday makes Android a prime candidate for online machine learning.

### v4

@@ 8.2 Comparison of ensemble models
9 Discussion
| Study | Accuracy | F1-Score | TPR | |||
| Original | Ours | Original | Ours | Original | Ours | |
| Static analysis models | ||||||
| Permissions and API calls [Peiravian2013] | 0.969 | 0.961 | — | 0.959 | 0.948 | 0.961 |
| Normal/dangerous permissions [Wang2014] | 0.996 | 0.930 | 0.895 | 0.946 | 0.926 | 0.931 |
| Selected permissions [Rathore2021] | 0.940 | 0.949 | — | 0.959 | 0.930 | 0.951 |
| Model-selected permissions [Sahin2023] | — | 0.924 | 0.956 | 0.925 | — | 0.932 |
| API calls with deep learning [Ma2019] | — | 0.970 | 0.990 | 0.969 | 0.988 | 0.979 |
| Reduced API calls [Jung2018] | 0.999 | 0.762 | — | 0.799 | — | 0.948 |
| Selected API calls [Muzaffar2021] | 0.962 | 0.968 | 0.961 | 0.968 | 0.959 | 0.965 |
| Drebin feature set [Arp2014] | 0.940 | 0.956 | — | 0.895 | — | 0.961 |
| Opcodes [Kang2016] | — | 0.973 | 0.980 | 0.975 | — | 0.965 |
| Opcodes with 2D deep learning [Xiao2019a] | 0.930 | 0.913 | 0.940 | 0.916 | 0.944 | 0.921 |
| Opcodes with 1D deep learning [yeboah2022] | — | 0.936 | 0.970 | 0.935 | 0.970 | 0.923 |
| Dynamic analysis models | ||||||
| Selected system calls [Ananya2020] | 0.994 | 0.951 | — | 0.950 | — | 0.952 |
| System calls with deep learning [Malik2019] | 0.852 | 0.820 | 0.859 | 0.806 | 0.839 | 0.752 |
| System and API calls [Afonso2015] | 0.968 | 0.962 | 0.968 | 0.964 | 0.961 | 0.959 |
| TCP network features [Zulkifli2018] | 0.984 | 0.977 | — | 0.981 | 0.920 | 0.978 |
| TCP and HTTP features [Wang2016] | 0.997 | 0.973 | — | 0.977 | — | 0.973 |
| Hybrid analysis models | ||||||
| TCP features and permissions [Kandukuru2017] | 0.956 | 0.943 | — | 0.945 | 0.950 | 0.948 |
| Permissions and system calls [Kapratwar2017] | — | 0.968 | — | 0.946 | — | 0.969 |
| Ensemble models | ||||||
| Dynamic system + static API calls (ours) | 0.978 | 0.976 | 0.979 | |||
| Dynamic system + static API calls (ours) | 0.971 | 0.971 | 0.971 | |||
| System/API calls + permissions + network (ours) | 0.969 | 0.970 | 0.966 | |||
This study allowed us to take a broader perspective on the feature-based Android malware-detection literature and—arguably for the first time—see how previously published approaches compare and perform within a contemporary Android environment. Table 33 summarises the results. Perhaps most importantly, these show that feature-based approaches can still perform very well within a contemporary context, with the best models achieving F1 scores of around 98%. This exceeds many of the figures published in recent studies that use transformers and end-to-end learning, but without the added risks around overfitting, transparency, execution speed and computational costs that these approaches bring [Lones2021]. Whilst this does not imply that more modern approaches are unable to achieve better performance than traditional machine learning, it does question the sometimes implicit assumption that more recent models are necessarily better. This suggests that the Android malware detection community should approach developments in AI in a more nuanced manner, building on the strengths of feature-based approaches.
Although the best feature-based malware detection models still achieve levels of performance inline with those originally published, we also observed that a number of approaches performed significantly less well when reimplemented and evaluated on a large contemporary dataset. This is despite us performing hyperparameter optimisation and often broadening the scope of studies, both of which might be expected to produce better models. Some of this difference may be due to distribution shifts in malware and benign software since the original studies were published, plus the increasing complexity of the Android environment. However, poor experimental practice is also likely to be a contributing factor. Whilst the aim here is not to diminish the value of the individual studies, we have observed many instances of common errors of practice [Lones2021] which may have led to over-confident measures of performance. A typical example of this is performing feature selection prior to splitting, which potentially leads to information leakage from the test data. Within this context, results of several of the studies dropped from near-perfect classification scores to relatively low scores.
The falls in accuracy are particularly pronounced. Table 33 shows that for those studies which did originally report accuracy figures, 77% (i.e. 10 studies) of them performed less well in our reimplementation, with the mean dropping from 0.961 to 0.928. This is a considerable drop, and the fact that this particularly affects accuracy may suggest that the figures reported in the original studies were unreliable due to their use of imbalanced datasets. For F1 scores, although 78% (7) performed less well after reimplementation, the difference in means is less pronounced, dropping from 0.947 to 0.934. However, fewer studies published F1 figures, so this distribution may be skewed towards those studies which followed better practice generally, including the reporting of a suitable selection of metrics. Nevertheless, regardless of the underlying cause, the reduction in performance shows the importance of periodically reimplementing and reevaluating past studies.
Another important observation is the challenge caused by the growing complexity of Android. For instance, the number of API calls has grown by two orders of magnitude since early work on ML-based Android malware detection. This means the use of feature selection is now very important in order to reduce feature numbers down to manageable levels. Within this context, an important observation is that feature selection does not just reduce the computational resources needed for training; it also, in most cases, increases the accuracy of the resulting models. Most notably, we have shown that reducing the number of API calls by 95% leads to higher levels of malware discrimination. In general, we found mutual information and chi-square to be the most effective feature selectors. This is a useful observation, since these model-agnostic approaches are relatively cheap, which is important when working with large feature sets. Model-based selection, which attempts to fit features to a particular downstream model, is more expensive, but was not found to yield better results.
@@ 9 Discussion
Some of the observations around feature utility are surprising. For instance, our results suggest that, within the scope of our evaluation, there is generally little advantage to using dynamic analysis techniques, and that models with similar accuracy can be built using relatively simple static features. The only scenario where dynamic analysis seemed to bring a marginal advantage was when we built models based around network activity. However, this is expensive to implement in practice, since collection of network activity requires dynamic analysis at both training and inference time, and relies on features which are both difficult to collect and potentially brittle.
@@
The Drebin feature set has historically played an important role in ML-based malware detection studies. Although this feature set still performs reasonably well, our results suggest that it is no longer the best feature set to use, and is difficult to deploy due to the large number of features it entails when used within a contemporary Android environment. Unlike other feature sets we studied, it does not benefit from feature selection.
@@
Beyond features, we also found that model choice is important, and random forests were surprisingly consistent in being the best performing models across the majority of experiments and feature sets. SVMs also performed well in general, but were more expensive to train than random forests. Naïve Bayes models generally performed poorly, suggesting that, for malware detection, interactions between features are important. A somewhat surprising observation, given their increasing prevalence within ML more generally, is that deep neural network models did not perform particularly well. However, this may be due to the use of pre-extracted features in the majority of malware detection studies.
10 Limitations and Future Work
@@
Our aim in this study was to sample and reimplement representative feature-based ML studies from the literature. It is not feasible to reimplement all existing studies, and there is inevitably some sampling bias within our approach. In particular, we focus on studies which used similar underlying features, since this enables a more meaningful basis for comparison across a standardised data set. This approach was motivated by our earlier review [Muzaffar2022] of the most common ML-based modelling approaches in Android malware detection. However, it should be noted that there are other studies which have used broader information sources to construct features. For example, Hou et al. [Hou2017], in their HinDroid framework, consider the relationships between applications and APIs. Similarly, Ye et al. [ye2019out] utilized features that incorporated various forms of metadata from apps and devices, including IMEI numbers. In future work, it would be interesting to implement and compare studies that incorporate these more heterogeneous features.
Although beyond the scope of this study, there are other characteristics of malware which can be important in practice. This includes the use of code obfuscation and packing methods, which are both sometimes used to evade detection. By assessing the detection of malware in bulk, it is unclear from our results how well different modelling approaches deal with this, and it would also be an interesting topic for a future study. Another important dimension within antimalware research is how to deal with the changing nature of malware over time. Whilst our study gives some insight into the robustness of particular modelling approaches in this regard, this was not a focus of this work, and we point interested readers to our recent work on using active learning as a means for responding to changes in the distribution of Android malware over time [actdroid].
Finally, it is likely that recent end-to-end learning and transformer-based approaches might suffer from the same issues around ML practice observed in feature-based studies. For this reason, it would also be useful to reimplement some of these more recent studies within a level playing field. However, the reimplementation of these models is beyond the scope of the framework developed for this current study.
11 Conclusions
This work revisited and reimplemented previous studies in feature-based ML for Android malware detection. The 18 foundational studies that we considered were originally published over a 10 year time period, from 2013–2023, during which both the Android operating system and the malware ecosystem evolved significantly. One of our aims aim was to determine how well these approaches work within a contemporary Android environment. Our results suggest they still hold up well, with the most effective models achieving accuracies and F1 scores of around 98%. This is more or less in line with the published performance of more recent approaches based around transformers and end-to-end learning.
However, overfitting is easily achieved when good practice is not followed. In this respect, it is concerning that the performance of most of the anti-malware models following reimplementation was less—and sometimes considerably less—than the figures originally published. This is likely to be at least partially attributable to issues of ML practice. We noted poor reporting standards (including the inappropriate use of metrics and the lack of statistical rigour) as a particularly common problem, but there are many other aspects of practice that can lead to over-confident results. We believe there is a need for better awareness of good ML practice in order to improve the replicability and robustness of research in this area.
@@ 11 Conclusions
A key observation was that features extracted using static analysis of Android APKs were often more effective than those extracted using dynamic analysis of running applications. This is important to know, since static analysis is much cheaper than dynamic analysis. Within this context, we found that simple representations of API usage led to very effective models, with accuracies of up to 96.8%. The exception to this is when features extracted from network traffic are used to build models. In our study, it is these models that provided the highest levels of malware discrimination, achieving accuracies of up to 97.7%. However, they are expensive to train and use, and rely on features that are difficult to collect and potentially brittle. To address this, we also showed that ensemble models built from non-network features can achieve a similar level of discrimination (97.8%), offering a more practical alternative for real world deployment.
Other notable findings include the importance of using feature selection algorithms when dealing with the complexity of modern Android systems, the relatively poor performance of the historically-important Drebin feature set, and the high accuracy of random forest models in comparison to other ML models, including various deep learning models. This latter point is especially relevant in the face of current trends in anti-malware research, which are exploring increasingly complex ML models, including LLMs. Given that relatively simple models can reach very high accuracies, this questions the need to use slower more resource-intensive models that are in general more susceptible to overfitting and to adversarial attacks.
Appendix A DroidDissector feature extraction tool
This appendix provides full details of the automated analysis tool, DroidDissector [muzaffar2023droiddissector], which we used to extract static and dynamic features from Android applications. It comprises two sub-systems: one for extracting static features, and another for extracting dynamic features.
A.1 Static Analysis Sub-System
APK files are first reverse-engineered and decompressed using APKtool [APKTool] to decode the manifest file and smali files from the APK package. Permissions used by the application, the application package name and the different application components — activities, services, broadcast receivers, and content providers — are stated in the Android manifest file. APKtool decodes the dex (binary code) files into readable Java-like smali files. APKTool creates a directory for each APK file. The static analysis tool then analyzes the files produced by the APKTool to retrieve features. For each application, the features extracted by our static analysis tool include:
- Hardware Components:
- Requested Permissions:
- App Components:
- Filtered Intents:
- Used Permissions:
- Network Addresses:
- API Calls v30:
- Restricted API Calls:
List of restricted API calls defined by Drebin [Arp2014]. We built this feature set by analyzing the smali files extracted from the APK.
- Suspicious API Calls:
List of suspicious API calls defined by Drebin [Arp2014]. We built this feature set by analyzing the smali files extracted from the APK.
- API Call Graphs:
- Opcodes:
A.2 Dynamic Analysis Sub-System
The dynamic analysis sub-system runs on the readily-available Android emulators provided by Android SDK [android_sdk]. The virtual images in the emulators also allow root access, which is essential for dynamic analysis. Here we describe the environment required to run the tool and the features extracted by the tool.
A.3 Environment
Python 3: All the dynamic modules require Python to run.
Android Emulator: Our tool uses the emulator provided by Android SDK as the virtual environment to run the analysis. The emulator should be on the system path. We have tested the tool on Android versions up to and including the current release (version 15).
A.4 Feature Extraction
- System Calls:
- Network Traffic:
Our tool uses Android emulator’s tcpdump to collect the complete network traffic of the application. It stores the resulting data in “pcap” files, later used to analyze and extract features. Wireshark [Wireshark] can process this file to analyze the application’s network traffic. It also saves a separate network traffic file for every application.
- Android Logcat:
- API Calls:
The tool requires hooking of API calls beforehand to monitor execution of API calls during the run-time of an application. However, it is not computationally feasible to monitor the 130,000+ API calls present in the Android SDK. Therefore, the tool carries out the following steps: before starting the dynamic analysis, it checks if a corresponding static analysis report is present for the application. In addition to hooking the classic calls used in the literature of dynamic analysis tools [mobsf], we also hooked application-specific API calls. If present, it hooks the API calls extracted in the static analysis report, or else, it performs a static analysis on the application using the static analysis tool. The dynamic analysis tool then hooks the API calls to monitor during the execution of the application using Frida. The tool saves a separate file with all the APIs called during execution for further analysis. As the process may involve static analysis, the API call module is optional in the dynamic analysis tool. By default, the tool monitors a pre-defined list of API calls; however to monitor the API calls mined from an application’s static analysis, the static analysis module is required.